OpenStudy (anonymous):

Missy is organizing the seating arrangements for a concert in her local park. The first row's seats are modeled by the function f(x) = 10. To give everyone a good view, each row must have two more seats than the row before it. The number of additional seats a row has is modeled by the function s(x) = 2(x - 1). Explain to Missy how to create an equation to predict the number of seats in any row. Describe how to determine the number of seats in the 20th row. Use complete sentences.

OpenStudy (anonymous):

@Abhisar @Luigi0210 or @paki please hep me ): so confused

pooja195 (pooja195):

Do you know the equation?

OpenStudy (anonymous):

no ): completely lost its like s(x)=(1-1) something like that

pooja195 (pooja195):

ok so fx=10 thats the first part of your equation \[x=10\]

OpenStudy (anonymous):

like I don't know what numbers to plug in and what number to multiply and subtract to get my answer in a non confusing way

pooja195 (pooja195):

oh im confused i go the equation but i dont know if i got it right because i got a = 10 + (19)2 because fx=10 and then since shes in the 20th row you plug in \[s(x)=2(20-1)=~x=10+19(2)\]

pooja195 (pooja195):

im pretty sure thats correct and now you just have to solve

pooja195 (pooja195):

Do you get that?

OpenStudy (anonymous):

Pooja why do you hate me \[D:\]

pooja195 (pooja195):

@izthewiz7 are you understanding what i said?

OpenStudy (anonymous):

I honestly don't understand this..... this is really depressing me, I've been on the same freaking question )):

pooja195 (pooja195):

ok lets do this step by step \[x=10\] you know this info already \[s(x) = 2(x - 1)\] this is the formula given to find the number of additional seats \[s(x)=2(20−1)\] we plug in 20 because that the only info you have on the seats disregarding the numbers 2 and all of that \[s(x)=2(20−1)= x=10+19(2)\] combine the solution of the equation with x=10 \[x=10+19(2)\] solve the equation to get answer

OpenStudy (anonymous):

Ik how you feel, Ive been on open study for two days and I know right now you just want the answer. You don't care about how you get it. I know you can get stuck on one question for a long time, but I assure you. Once you earn the answer and you get the feeling of accomplishment it Is so worth it. :) Just hang in there

pooja195 (pooja195):

@izthewiz7 do you get it now?

OpenStudy (anonymous):

@pooja195 48? and @EinsteinPart2 yeah, I've been on this problem for 2 hours, I just need to understand because I kave another question like this /:

pooja195 (pooja195):

No its not 48 you did something wrong

OpenStudy (anonymous):

I give up /: thanks though, I really appreciate it..

pooja195 (pooja195):

no dont give up you just have to solve that last equation :D come on you can do it :D

OpenStudy (anonymous):

I get so confused at the part of when you got to s(x)= 2(20-1) = x = 10 + 19(2)

OpenStudy (anonymous):